# Constrained Entropic Unlearning: A Primal-Dual Framework for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) deployed in real-world settings increasingly face the need to unlearn sensitive, outdated, or proprietary information. Existing unlearning methods typically formulate forgetting and retention as a regularized trade-off, combining both objectives into a single scalarized loss. This often leads to unstable optimization and degraded performance on retained data, especially under aggressive forgetting. We propose a new formulation of LLM unlearning as a constrained optimization problem: forgetting is enforced via a novel logit-margin flattening loss that explicitly drives the output distribution toward uniformity on a designated forget set, while retention is preserved through a hard constraint on a separate retain set. Compared to entropy-based objectives, our loss is softmax-free, numerically stable, and maintains non-vanishing gradients, enabling more efficient and robust optimization. We solve the constrained problem using a scalable primal-dual algorithm that exposes the trade-off between forgetting and retention through the dynamics of the dual variable. Evaluations on the TOFU and MUSE benchmarks across diverse LLM architectures demonstrate that our approach consistently matches or exceeds state-of-the-art baselines, effectively removing targeted information while preserving downstream utility.
@@ 1 Introduction
Machine unlearning, initially introduced by Cao and Yang [5], poses a fundamental challenge: how can one remove the impact of a small subset of data without retraining the model from scratch? For LLMs, full retraining is prohibitively expensive, especially as models grow in size. Additionally, LLMs must frequently forget information to comply with regulatory mandates (e.g., the “right to be forgotten”  [43]), avoid generating harmful content [41], prevent the leakage of private data [40], or eliminate reliance on copyrighted materials [9].
@@
This has motivated recent algorithmic efforts to approximate unlearning via lightweight fine-tuning techniques, most notably gradient ascent over the forget set [41, 9, 25]. While such methods can suppress the model’s ability to recall or generate content related to the undesired data, they often do so at a steep cost: they degrade model performance on unrelated, retained data. This degradation is particularly problematic when the forget set is much smaller than the retained dataset, as is typically the case in practice. As a result, recovering the lost utility of the model becomes extremely difficult and costly
Prior work has proposed theoretically principled formulations of unlearning as a regularized optimization problem, seeking model parameters that (i) degrade performance on a forget dataset to remove sensitive or undesirable knowledge, and (ii) preserve performance on a retain dataset to maintain utility. However, this framework encounters three key obstacles.
First, regularized formulations treat unlearning as a soft trade-off between forgetting and retention. In doing so, they often drive the forget loss to vanishingly small values (almost zero e.g.,  [4]) in an effort to eliminate all traces of the undesired data. However, this aggressive forgetting is not only computationally inefficient, but also often counterproductive, leading to a sharp degradation in the utility of the model in unrelated inputs.
Second, at the beginning of unlearning, the initial utility of the model is already high, i.e., the retention loss is small. The regularized formulation attempts to decrease the retention loss at the same time as the forgetting loss, but this is unnecessary: Simply keeping the retention loss at its initial level is enough to maintain the original utility of the model. Trying to reduce the retention loss like the forgetting loss can lead to (i) overfitting and bias toward the retained set, and (ii) difficulty in sufficiently lowering the forgetting loss, since the problem becomes a multi-objective optimization with conflicting gradient directions, and as a result, both forgetting and retention losses decrease poorly.
Third, conventional loss functions such as cross-entropy are poorly behaved when repurposed for forgetting. Cross-entropy is unbounded above; thus, maximizing it results in unstable gradients, optimization divergence, and a lack of interpretability. These behaviors hinder convergence and offer little guidance on when forgetting is sufficient.
Our Contributions. (1) We propose a conceptual shift: unlearning should be cast as a constrained optimization problem, where the goal is to forget designated data while explicitly preserving utility on retained data via a formal constraint. This avoids the need for delicate loss balancing and provides strong performance guarantees. (2) We introduce a novel logit-margin flattening loss that encourages uniform output distributions on the forget set, offering a stable, softmax-free alternative to entropy maximization. This loss is convex, bounded, and provides non-vanishing gradients–making it well-suited for large-scale optimization. (3) We develop a scalable primal-dual algorithm that enforces the retention constraint and captures the trade-off between forgetting and utility through the dynamics of the dual variable. The algorithm supports warm starts and dynamic updates, ensuring efficiency even at LLM scale. (4) Finally, we evaluate our method on the TOFU and MUSE benchmarks using both standard metrics and a new LLM-based judge that assesses behavioral divergence.
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.7Forget Success (higher is better)Model Utility (higher is better)TARGETRETRAINEDGRAD-DIFFDPONPOSIMNPORMUPDU (OURS)
Figure 1: Comparison of different methods on the TOFU dataset on Llama 3.2 3B: Model Utility vs. forget Success; see Section 4 for explanation of the metrics.
Figure 2: Samples from responses of different methods to same prompt in forgetting and retaining evaluations on the MUSE (Books)
@@ 1.1 Related Works
All existing methods prioritize unlearning performance on the forget set. Overoptimization towards minimizing forget loss can disproportionately degrade model utility, which is challenging to recover given the extensive nature of retain sets. As shown in Figure 2, even retraining the model on the retain data set does not achieve an extremely low loss on the forget set. This suggests that pushing the forget loss lower leads to overforgetting, which makes it easier for later attacks to relearn the forgotten data, since forgetting is not uniform [14]. Another benefit of our method is that it enables a dynamic approach towards handling conflicting gradients. That is, when optimizing two or more losses [41], it is likely that gradients will be conflicting and finding a shared descent direction can be extremely resource intensive for large scale LLMs [27]. However, the primal dual algorithm naturally adjusts the linear combination such that if the retain loss is below our desired threshold, its conflicting dynamic with the forgetting loss could be completely ignored. The dynamicaly changing $\lambda$ would promptly shift attention back to the retention loss if the loss exceeds the user-defined threshold. This approach allows the model to unlearn the forget set while remaining as close as possible to the original model in utility.
@@ 3 Constrained Entropic Unlearning
The linear scalarization formulation in (1) suffers from a limitation: if the forget–set $\mathcal{D}_{\text{fgt}}$ and retain-set $\mathcal{D}_{\text{rtn}}$ overlap (statistically or semantically), blindly reducing $\mathcal{L}_{\text{fgt}}$ can catastrophically increase $\mathcal{L}_{\text{rtn}}$. To balance this trade-off, the scalarization weight $\lambda$ must be carefully tuned for each instance. However, even such a dynamic approach provides no explicit control over the degradation on the retention set. In particular, small values of $\lambda$ may lead to incomplete forgetting, while large values can overly compromise retention.
@@ 3.1 Lagrangian Relaxation and the Dual Problem
[…]
@@
Here, the search space is restricted to $\Pi_{\theta}=\{\pi_{\theta}\mid\theta\in\Theta\}\subseteq\Pi$. While strong duality may not strictly hold in this finite-dimensional, nonconvex setting, modern models are typically sufficiently expressive to approximate the infinite-dimensional problem well [10].
@@ 3.2 Proposed Method: Primal-dual with Warm Start
While dual ascent offers strong theoretical guarantees, it typically involves a costly inner-loop optimization to fully minimize the Lagrangian at each step. We propose an efficient variant that performs a single warm-started dual ascent step, followed by lightweight primal-dual updates. The initial iteration fully minimizes $\mathcal{L}({\pi_{\theta}},\lambda_{0})$ with respect to $\theta$. Subsequent iterations alternate between one gradient descent step on $\theta$ and one dual ascent step on $\lambda$, reducing computation through single-step updates while retaining the advantages of dual ascent initialization. This method is detailed in Algorithm 1.
@@
8:         $\theta\leftarrow\theta-\eta_{\theta}\nabla_{\theta}\mathcal{L}(\theta,\lambda)$triangleright $\nabla_{\theta}\mathcal{L}(\theta,\lambda)=\nabla_{\theta}\ell_{\mathrm{f}}(\theta)+\lambda\nabla_{\theta}\ell_{\mathrm{r}}(\theta)$
9:         if $t>T_{\mathrm{w}}$then $\triangleright$ Warm-start; Solve the primal problem for a fixed $\lambda$ until epoch $T_{\mathrm{w}}$
10:              $\lambda\leftarrow\left[\lambda+\eta_{\lambda}\left(\ell_{\mathrm{r}}(\theta)-\varepsilon\right)\right]_{+}$triangleright Dual update and project onto $\mathbb{R}_{\geq 0}$
@@
While our framework is compatible with a broad class of loss functions proposed in prior unlearning literature, we will focus on specific instantiations of $\mathcal{L}_{\text{fgt}}$ and $\mathcal{L}_{\text{rtn}}$, deferring extensions to alternative losses to the Supplementary Material.
@@ 3.4 Logit Flattening for Efficient Forgetting
However, the CE loss is unbounded above, and directly maximizing it during unlearning often leads to gradient explosion and catastrophic collapse. Notably, CE minimization during pretraining serves as an upper bound surrogate for the 0-1 classification loss. Reversing this objective, by maximizing CE, invalidates this surrogate relationship
@@
While effective, entropy-based losses suffer from vanishing gradients as the output distribution approaches uniformity, which can slow down convergence and cause instability in later stages of training. Moreover, they require the log-softmax computation, which can be numerically sensitive and costly in large vocabulary settings.
@@ 4 Experiments
We evaluated our unlearning methodology on two established benchmarks: TOFU and MUSE [25, 33, 8]. The TOFU dataset consists of 200 diverse fictional author profiles, each containing 20 question-answer pairs. A designated subset of these profiles, known as the forget set, serves as the target for unlearning. In the main experiments, we choose to forget the subset Forget10 and defer Forget05 and Forget01 to the Supplementary Material. The MUSE benchmark focuses on unlearning in two real-world contexts: Books and News. Due to space constraints, we consider the News subset in the main experiments and defer Books to the Supplementary Material. The News subset is a collection of BBC news articles collected after August 2023.
@@ Methods:
We compare our method, Primal-Dual Unlearning (PDU), against several baselines. The first is the target model that has been trained on $\mathcal{D}_{\text{rtn}}\cup\mathcal{D}_{\text{fgt}}$. Second, we consider an ideal model that has only been trained on $\mathcal{D}_{\text{rtn}}$. Next, we turn to several established methods: GradDiff [41], DPO [25], NPO [45], SimNPO [11], and RMU [21]. For all the baselines, we use $\lambda=1$ as the trade-off parameter.
@@
  • Forget Success: We define this metric as the harmonic mean over $1-$ the likelihood on $\mathcal{D}_{\text{fgt}}$, $1-$ the ROUGE score on $\mathcal{D}_{\text{fgt}}$, and the truth ratio on $\mathcal{D}_{\text{fgt}}$. For metric definitions see [25].
@@ Results:
The results of our experiments are reported in Table 1 and Table 2. For unlearning, it is critical to take into account the involved metrics at the same time and to not judge based on single metrics. That is, a model can have a perfect forgetting score but at the cost of catastrophic forgetting severely hampering the models utility.
@@
We can see in Table 1 that our method consistently outperforms all other methods across various scales and models by achieving the highest forget successes whilst retaining the best or second best model utilities. We see similar exceptional performance from our methodology on the LLM judged metrics, except for the Fluency metric. Upon further examination, it becomes clear that this is an artifact of the success of the unlearning algorithm. That is, on the forget set the model’s knowledge has been purged and the model abstains from making coherent predictions. Importantly, it should be noted that the model fluency on the other tasks is unaffected. Due to space constraints, we defer the detailed LLM judge statistics to the Supplementary Material.
@@
We see that for the larger 7B and 8B models, GradDiff has a forget success of 00 but an LLM judged forget score of $10$. Studying the generations of the models, we see that the models unlearned via GradDiff abstain from producing any text when prompted with prompts from $\mathcal{D}_{\text{fgt}}$. As such the truth ratio on $\mathcal{D}_{\text{fgt}}$ is essentially 00 and yields a 00 harmonic mean for the forget success. Due to this behaviour, the judge gives a complete forgetting score to the model. However, unlike our method, we see that GradDiff suffers from this artifact in its utility and also the other LLM judged metrics.
We further see that our method performs competitively on the more complex task of MUSE-News per Table 2. That is, our method provides a viable pareto point that provides unique retain and forget ROUGE scores.
@@
Table 1 and Table 2 further point to an important observation: the traditional metrics used for assessing task success, i.e., metrics such as model utility and forget success, are generally indicative of real success, as outlined by the correlation that we see with the LLM judged metrics. Without the LLM judged metrics, it was not clear if metrics such as the likelihood of generating the prompt-response pair $(x,y)\in\mathcal{D}_{\text{fgt}}$ or the ROUGE score would be real indicators of the successful unlearning. The LLM judged metrics show that this is generally the case and classical metrics are still useful indicators of a model’s capabilities.

### v2

@@ Abstract
Large Language Models (LLMs) deployed in real-world settings increasingly face the need to unlearn sensitive, outdated, or proprietary information. Existing unlearning methods typically formulate forgetting and retention as a regularized trade-off, combining both objectives into a single scalarized loss. This often leads to unstable optimization and degraded performance on retained data, especially under aggressive forgetting. We propose a new formulation of LLM unlearning as a constrained optimization problem: forgetting is enforced via a novel logit-margin flattening loss that explicitly drives the output distribution toward uniformity on a designated forget set, while retention is preserved through a hard constraint on a separate retain set. Compared to entropy-based objectives, our loss is softmax-free, numerically stable, and maintains non-vanishing gradients, enabling more efficient and robust optimization. We solve the constrained problem using a scalable primal-dual algorithm that exposes the trade-off between forgetting and retention through the dynamics of the dual variable, all without any extra computational overhead. Evaluations on the TOFU and MUSE benchmarks across diverse LLM architectures demonstrate that our approach consistently matches or exceeds state-of-the-art baselines, effectively removing targeted information while preserving downstream utility.
@@ 1 Introduction
Machine unlearning, initially introduced by Cao and Yang [7], asks a fundamental question: how can one remove the impact of a small subset of data without retraining the model from scratch? For LLMs, full retraining is prohibitively expensive, especially as models grow in size. Additionally, LLMs must frequently forget information to comply with regulatory mandates (e.g., the “right to be forgotten”  [52]), avoid generating harmful content [50], prevent the leakage of private data [49], or eliminate reliance on copyrighted materials [13].
@@
This has motivated recent algorithmic efforts to approximate unlearning via fine-tuning techniques, most notably gradient ascent over the forget set [50, 13, 29]. While such methods can suppress the model’s ability to recall or generate content related to the undesired data, they often do so at a steep cost: they degrade model performance on unrelated, retained data. This degradation becomes especially severe when the forget set is small relative to the retained corpus, a common real-world setting, making the recovery of model utility both difficult and costly.
In principle, unlearning can be formulated as a multi-objective optimization problem, as explored in recent works [5, 32]. This formulation naturally captures the trade-off between minimizing the forget loss and preserving performance on the retain set. Dedicated multi-objective gradient methods can, in theory, achieve a balanced Pareto-optimal solution. However, these algorithms typically require multiple gradient evaluations per iteration, making them computationally impractical for large-scale models, particularly commercial LLMs with tens or even hundreds of billions of parameters.
The prevailing approach to unlearning, regularizing the forget loss via an additional penalty that discourages model degradation [25, 50], can be interpreted as a linear scalarization of the underlying multi-objective problem. While conceptually straightforward, this approach suffers from two fundamental limitations. First, from a theoretical standpoint, linear scalarization cannot recover the entire Pareto frontier in multi-objective optimization [4, 3, 10]. As a result, regularized unlearning explores only a limited subset of feasible trade-offs, potentially excluding solutions that offer more balanced performance between forgetting and retention. Second, from a practical perspective, the regularization coefficient is often opaque and requires extensive, task-specific tuning, which complicates deployment and hinders reproducibility.
In this work, we take a step back and revisit the foundational multi-objective formulation of unlearning. Rather than relying on computationally intensive multi-gradient algorithms or heuristic regularization schemes, we cast LLM unlearning as an $\varepsilon$-constrained optimization problem [30]. In this formulation, one objective is transformed into an explicit constraint with an interpretable threshold $\varepsilon$, offering direct control over the trade-off between forgetting and utility preservation. This perspective simultaneously mitigates the theoretical limitations of linear scalarization and the computational overhead of multi-objective methods, while yielding a more principled and scalable framework for large-scale unlearning. We summarize our main contributions as follows:
  • We formulate LLM unlearning as a constrained optimization problem, where the objective is to erase designated knowledge while explicitly enforcing utility preservation on retained data through an explicit constraint. This formulation removes the need for delicate loss balancing and provides clear theoretical guarantees.
  • We propose a logit-margin flattening loss that promotes uniform model outputs on the forget set, serving as a stable, softmax-free alternative to entropy maximization. The loss is convex, bounded, and yields non-vanishing gradients, making it suitable for large-scale optimization.
  • We design a scalable primal-dual algorithm that enforces the retention constraint and naturally captures the forgetting-utility trade-off through the dynamics of the dual variable. The method supports warm starts and dynamic updates, achieving efficiency at LLM scale with no additional gradient computations.
  • We validate our approach on the TOFU and MUSE benchmarks using both standard metrics and a novel LLM-based judge to evaluate behavioral divergence.
@@ 1.1 Related Works
00.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.91100.10.10.20.20.30.30.40.40.50.50.60.60.70.7Forget Success (higher is better)Model Utility (higher is better)TARGETRETRAINEDGRAD-DIFFDPONPOSIMNPORMUPDU (OURS)
Figure 1: (left) Comparison of different methods on the TOFU dataset on Llama 3.2 3B: Model Utility vs. forget Success; see Section˜4 for explanation of the metrics. (right) Overview of our methodology and contributions. Unlearning is cast as a constrained optimization, then solved using primal dual optimization. We use our novel logit-flattening loss for the forgetting task. When the retention loss is larger than the pre-specified threshold, dual updates increase the value of $\lambda$, increasing the effect of the retention loss. When the retention loss is in the desired range, the dual updates reduce $\lambda$ so that the optimization can tackle the forget loss more effectively.
@@
Existing methods predominantly emphasize minimizing the loss on the forget set. Such over-optimization often causes disproportionate deterioration of general utility on the retain set, and the resulting performance gap is difficult to recover given the breadth and complexity of the retained data distribution. As shown in Figure 1, even retraining the model on the retain data set does not achieve an extremely low loss on the forget set. This suggests that pushing the forget loss lower leads to overforgetting, which makes it easier for later attacks to relearn the forgotten data, since forgetting is not uniform [18]. Another benefit of our method is that it enables a dynamic approach towards handling conflicting gradients. That is, when optimizing two or more losses [50], it is likely that gradients will be conflicting and finding a shared descent direction can be extremely resource intensive for large scale LLMs [32]. However, the primal dual algorithm naturally adjusts the linear combination such that if the retention loss is below our desired threshold, its conflicting dynamic with the forgetting loss could be completely ignored. The dynamically changing regularization would promptly shift attention back to the retention loss if the loss exceeds the user-defined threshold. This approach allows the model to unlearn the forget set while remaining as close as possible to the original model in utility.
@@ 3 Constrained Entropic Unlearning
The linear scalarization formulation in (1) suffers from a limitation: if the forget‐-set $\mathcal{D}_{\text{fgt}}$ and retain‐set $\mathcal{D}_{\text{rtn}}$ overlap (statistically or semantically), reducing $\mathcal{L}_{\text{fgt}}$ can increase $\mathcal{L}_{\text{rtn}}$. To balance this trade-off, the scalarization weight $\lambda$ must be carefully tuned for each instance. However, even such a dynamic approach provides no explicit control over the degradation on the retention set. In particular, small values of $\lambda$ may lead to incomplete forgetting, while large values can overly compromise retention.
@@ 3.1 Lagrangian Relaxation and the Dual Problem
$\text{(Primal)}\quad\min_{\pi\in\Pi}\max_{\lambda\geq 0}\mathcal{L}(\pi,\lambda),\qquad\text{(Dual)}\quad\max_{\lambda\geq 0}\min_{\pi\in\Pi}\mathcal{L}(\pi,\lambda).$
@@
Here, the search space is restricted to $\Pi_{\theta}=\{\pi_{\theta}\mid\theta\in\Theta\}\subseteq\Pi$. While strong duality may not hold in this finite-dimensional, nonconvex setting, modern models are typically sufficiently expressive to approximate the infinite-dimensional problem well [14].
@@ 3.2 Proposed Method: Primal-dual with Warm Start
While dual ascent offers strong theoretical guarantees, it typically involves a costly inner-loop optimization to fully minimize the Lagrangian at each step. We propose an efficient variant that performs a single warm-started dual ascent step, followed by primal-dual updates. The initial iteration fully minimizes $\mathcal{L}({\pi_{\theta}},\lambda_{0})$ with respect to $\theta$. Subsequent iterations alternate between one gradient descent step on $\theta$ and one dual ascent step on $\lambda$, reducing computation through single-step updates while retaining the advantages of dual ascent initialization. This method is detailed in Algorithm˜1. Importantly, the proposed implementation in Algorithm˜1 incurs no extra computational overhead compared to linear regularization methods.
@@
8:   $\theta\leftarrow\theta-\eta_{\theta}\nabla_{\theta}\mathcal{L}(\theta,\lambda)$ $\triangleright$ $\nabla_{\theta}\mathcal{L}(\theta,\lambda)=\nabla_{\theta}\ell_{\mathrm{f}}(\theta)+\lambda\nabla_{\theta}\ell_{\mathrm{r}}(\theta)$
9:   if $t>T_{\mathrm{w}}$. then $\triangleright$ Warm-start; Solve the primal problem for a fixed $\lambda$ until epoch $T_{\mathrm{w}}$
10:     $\lambda\leftarrow\left[\lambda+\eta_{\lambda}\left(\ell_{\mathrm{r}}(\theta)-\varepsilon\right)\right]_{+}$ $\triangleright$ Dual update and project onto $\mathbb{R}_{\geq 0}$
@@
While our framework is compatible with a broad class of loss functions proposed in prior unlearning literature, we will focus on specific instantiations of $\mathcal{L}_{\text{fgt}}$ and $\mathcal{L}_{\text{rtn}}$. We establish these next.
@@ 3.4 Logit Flattening for Efficient Forgetting
However, the CE loss is unbounded above, and directly maximizing it during unlearning often leads to gradient explosion and catastrophic collapse. Notably, CE minimization during pretraining serves as an upper bound surrogate for the 0-1 classification loss. Reversing this objective, by maximizing the CE, invalidates this surrogate relationship and forfeits its theoretical justification.
@@
While effective, entropy-based objectives exhibit vanishing gradients, which slows convergence and destabilizes late-stage training. They also require the log-softmax over the full vocabulary, which is numerically sensitive and computationally heavy for large $V$.
@@ 4 Experiments
We evaluated our unlearning methodology on two established benchmarks: TOFU and MUSE [29, 39, 12]. The TOFU dataset consists of 200 diverse fictional author profiles, each containing 20 question-answer pairs. A designated subset of these profiles, known as the forget set, serves as the target for unlearning. In the main experiments, we choose to forget the subset Forget10 and defer Forget05 and Forget01 to the Supplementary Material. The MUSE benchmark focuses on unlearning in two real-world contexts: Books and News. While the TOFU dataset tests unlearning under a well-controlled setting, the MUSE benchmark presents a more challenging scenario with high overlap and imbalance between the forget and retain sets. The News subset is a collection of BBC news articles collected after August 2023. The Books subset presents an especially challenging scenario: unlearn copyrighted information from the Harry Potter book series, whilst retaining publicly available knowledge from Harry Potter Fan Wiki.
@@ Methods:
We compare our method, Primal-Dual Unlearning (PDU), against several baselines. The first is the target model that has been trained on $\mathcal{D}_{\text{rtn}}\cup\mathcal{D}_{\text{fgt}}$. Second, we consider an ideal model that has only been trained on $\mathcal{D}_{\text{rtn}}$. Next, we turn to several established methods: GradDiff [50], DPO [29], NPO [55], SimNPO [15], and RMU [25]. We utilize the OpenUnlearning GitHub repository for all the implementations. Moreover, our algorithm is implemented in this repository and made public at https://github.com/locuslab/open-unlearning.
@@
  • Forget Success: We define this metric as the harmonic mean over 11- the likelihood on $\mathcal{D}_{\text{fgt}}$, 11- the ROUGE score on $\mathcal{D}_{\text{fgt}}$, and the truth ratio on $\mathcal{D}_{\text{fgt}}$. For metric definitions see [29].
@@ Results:
The results of our experiments are reported in Table˜1 and Table˜2. When evaluating unlearning, it is critical to have a holistic view of the different metrics. That is, a successful unlearning is one that retains an acceptable level of model utility whilst forgetting the undesired data. For example, for the TOFU benchmark in Table˜1, an algorithm that has a very high model utility but a poor forget success has not been successful and has not forgotten the information. On the other hand, an algorithm that has a very desirable forget success but also has a significant reduction in model utility has degraded the model, making the model unappealing for production. To streamline comparisons, we provide an aggregating metric for success in Tables 1 and 2, which is simply the harmonic mean of the metrics.
@@
We can see in Table˜1 that our method consistently outperforms all other methods across various scales and models by achieving the highest forget successes whilst retaining the best or second best model utilities. We see similar exceptional performance from our methodology on the LLM judged metrics, except for the Fluency metric. Upon further examination, it becomes clear that this is an artifact of the success of the unlearning algorithm. That is, on the forget set the model’s knowledge has been purged and the model abstains from making coherent predictions. Importantly, it should be noted that the model fluency on the other tasks is unaffected. Due to space constraints, we defer the detailed LLM judge statistics to the Supplementary Material.
@@
We see that for the larger 7B and 8B models, GradDiff has a forget success of 0 but an LLM judged forget score of $10$. Studying the generations of the models, we see that the models unlearned via GradDiff abstain from producing any text when prompted with prompts from $\mathcal{D}_{\text{fgt}}$. As such the truth ratio on $\mathcal{D}_{\text{fgt}}$ is essentially 0 and yields a 0 harmonic mean for the forget success. Due to this behaviour, the judge gives a complete forgetting score to the model. However, unlike our method, we see that GradDiff suffers from this artifact in its utility and also the other LLM judged metrics.
We further see that our method performs competitively on the more complex tasks of the MUSE benchmark per Table˜2. For the MUSE Books task, we see that our method has achieved the most forgetting for both models from the methods that have not degenerated (GradDiff has significantly impaired the model and reduced its utility to near zero). For both the 7B and 13B models, our algorithm maintains high utility as observed via both the Retain ROUGE and the LLM-Judged Retain Score. For the MUSE News task, our method provides viable Pareto optimal points that provide unique retain and forget ROUGE scores.
@@
Table˜1 and Table˜2 further point to an important observation: the traditional metrics used for assessing task success, i.e., metrics such as model utility and forget success, are generally indicative of real success, as outlined by the correlation that we see with the LLM judged metrics. Without the LLM judged metrics, it was not clear if metrics such as the likelihood of generating the prompt-response pair $(x,y)\in\mathcal{D}_{\text{fgt}}$ or the ROUGE score would be real indicators of the successful unlearning. The LLM judged metrics show that this is generally the case and classical metrics are still useful indicators of a model’s capabilities.
Further Experiments and Evaluations
We provide a series of further experiments and evaluations which we defer to the Appendix due to space limitations. Section˜B.1 looks into providing a visualization of the results of Tables 1 and 2 through radar charts. Section˜B.3 establishes the details of our LLM-judge with samples. In Section˜B.6 we study longer unlearning using the different algorithms, study member inference attacks, exact memorization, and extraction strength, and conduct an ablation on single-turn jailbreak prompts to pique simple rephrasing attacks. See each corresponding section for details.